Please help, need help with finance worksheet problems to better understand, Thank you so much!!!!
What is the present value of the following annuities?
A. $2,500 a year for 10 years discounted back to the present at 7 percent

B. $70 a year for 3 years discounted back to the present at 3 percent
c. $280 a year for 7 years discounted back to the present at 6 percent
C. $500 a year for 10 years discounted back to the present at 10 percent

Answer by rfer(10417)

A) P=FV/(1+.07)^10
P=2500/(1.07)^10
P=2500/1.967151357
P=$1270.87
Here is (A) try the others on your own.
Good luck